OpenStudy (anonymous):

Hello there, I need help with this algebraic expressions problem. A= 3/2 B=5/2 c= 1/2 So 7-a+b-1-c. I keep getting 8 or some other wrong answer.

OpenStudy (anonymous):

\[7 - \frac{3}{2} + \frac{5}{2} - 1 - \frac{1}{2}\] since almost all the terms have a 2 in the denominator, it will be easiest to make the 7 and 1 have a 2 in the denominator as well, you do it like this \[\frac{14}{2} - \frac{3}{2} + \frac{5}{2} - \frac{2}{2} - \frac{1}{2}\] simplify \[\frac{14 - 3 + 5 - 2 - 1}{2} = \frac{13}{2}\]

OpenStudy (anonymous):

The answer is 7 1/2

OpenStudy (anonymous):

You're sure?? because what i got is correct and \[\frac{13}{2} = 6 \frac{1}{2}\]

OpenStudy (anonymous):

That is the answer in the book

OpenStudy (anonymous):

I'm.sorry. c is - 1/2

OpenStudy (anonymous):

Then either there is a typo or you copied the equation wrong because 7 - a + b - 1 - c = 13/2 with the values you have provided

OpenStudy (anonymous):

Yeah c is a negative 1/2 not positive the other variables are positive

OpenStudy (anonymous):

ok so \[\frac{14}{2} - \frac{3}{2} + \frac{5}{2} - \frac{2}{2} + \frac{1}{2} = \frac{14 - 3 + 5 -2 + 1}{2} = \frac{15}{2} = 7\frac{1}{2}\] thats better :)

OpenStudy (anonymous):

Wait for some reason I cannot see the rest of your evaluation after -2

OpenStudy (texaschic101):

7 - 3/2 + 5/2 - 1 - 1/2 (multiply the entire equation by 2) 14 - 3 + 5 - 3 - 1 19 - 7 = 12

OpenStudy (texaschic101):

wait...I made a mistake

OpenStudy (texaschic101):

7 - 3/2 + 5/2 - 1 - 1/2 (multiply by 2) 14 - 3 + 5 - 2 - 1 19 - 6 = 13

OpenStudy (anonymous):

Thank you Tyler but im still confused .why are you multiplying the whole thing by 2 ?

OpenStudy (anonymous):

You dont multiply the whole thing by 2

OpenStudy (anonymous):

you want all the terms to have a common denominator, since there were already 3 terms with a 2 in the denominator, it would be easiest to make the 7 have a 2 in the denominator and the 1 to have a 2 in the denominator.

OpenStudy (anonymous):

I did 7/1 -3/2 + 5/2 -1/1/ * -1/2

OpenStudy (anonymous):

right initially you have \[\frac{7}{1} - \frac{3}{2} + \frac{5}{2} - \frac{1}{1} + \frac{1}{2}\]

OpenStudy (anonymous):

Oh oh ohjhhhhhhhhhh. Now I get it I'm doing to much of theese math problems so I just experienced a brain fart haha

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

in order to add fractions, every fraction must have the same denominator

OpenStudy (anonymous):

You good now? lol

OpenStudy (anonymous):

Lol yeah I just need a break or something oh gosh

OpenStudy (anonymous):

Thanks for pointing that out and thank you Texas girl